# OpenSU3D: Open World 3D Scene Understanding using Foundation Models

### v1

@@ III-D Feature Fusion
However, this simple fusion schemes described in Eq. 4 and Eq. 5 is limited and offers lower performance on relative queries as highlighted later in ablation studies in Sec. V-A. Moreover, as also highlighted in findings from OpenMask3D [27], it was observed that multiscale crops add redundant context, with larger crops deteriorating overall object recall performance. To address these issues, we propose a modified multi-scale fusion scheme as shown in Eq. 6. Our proposed fusion scheme effectively down weighs the influence of larger crops by assigning weights based on cosine similarity to the best-fit crop.
$f_{\text{MS}}=\frac{1}{k}\sum_{i=1}^{k}\left(\frac{f_{1}\cdot f_{i}}{\max(\|f_{1}\|_{2}\cdot\|f_{i}\|_{2})}\right)\cdot f_{i}$ (6)
For the integration of multiview features, we draw parallels to the approach in ConceptFusion [11] which focuses on the fusion of 2D per-pixel features, we propose the direct incorporation of a global feature vector $f^{g}$, while synthesizing multiview feature vectors for each instance, defined as:
$f_{\text{MV}}=\frac{1}{m}\sum_{i=1}^{m}\left(f_{\text{MS}_{i}}+\left(\frac{f_{\text{MS}_{i}}\cdot f_{\text{G}_{i}}}{\max(\|f_{\text{MS}_{i}}\|_{2}\cdot\|f_{\text{G}_{i}}\|_{2})}\right)\cdot f_{\text{G}_{i}}\right)$ (7)
III-E Instance Retrieval & Segmentation
Given a map $\mathcal{M}$, open vocabulary 3D object search or 3D instance retrieval and segmentation operates in two stages. First, a query $\mathcal{K}$ is processed using the CLIP text encoder to obtain the feature vector $f_{\mathcal{K}}$. Second, a cosine similarity score is computed for all 3D instances {$S_{\text{score}}$}, and the segmentation mask of the 3D instance with maximum similarity score, $\mathcal{M}[\text{argmax}(\{S_{\text{score}}\})]$ is retrieved as the most likely response to the query $\mathcal{K}$.
III-F Spatial Reasoning
For queries involving complex spatial reasoning, the key idea is to leverage the long context window of Large Language Models like GPT-4 [6], to perform spatial reasoning based on coherent 3D representation and meta data such as mask labels, centroids, bounding boxes and description available for the constructed scene $\mathcal{M}$, using in-context learning. A simplified map :=$\mathcal{M}^{\prime}:=\mathcal{M}\setminus\{\mathcal{P},f_{MV}\}$ is passed along with system prompt designed using the prompting strategy defined as:
@@ III-F Spatial Reasoning
IV Experiment Setup
IV-A Datasets
Multiple scenes from the semi-synthetic dataset Replica [2] (room0, room1, room2, office0, office1, office2, office3, office4) and real-world ScanNet [1] (scene0000_00, scene0034_00, scene0164_03, scene0525_01, scene0549_00) were used for thorough qualitative and quantitative evaluation. Similar to previous studies [34, 11], a limited number of scenes were selected due to the extensive manual human evaluation.
IV-B Implementation Details
IV-B1 Models Utilized
@@
GroundedSAM [13] (a method based on RAM++ [35] (ram_plus_swin_large_14m), GroundingDINO [36] (groundingdino_swint_ogc), and SAM [4] (sam_vit_h_4b8939)) is employed for generating instance segmentation masks and bounding boxes. GPT-4V [6] (gpt-4-vision-preview, gpt-4-1106-preview) is used to create detailed captions & names for the instances and spatial reasoning. The CLIP encoder [3] (ViT-H-14 pre-trained on laion2b_s32b_b79k dataset) is used for instance feature vectors.
IV-B2 Hyper parameter Settings
@@
Same hyper parameters are used across datasets, determined by ablation studies on Replica [2] (Sec: V-A). We set $m=5$, selecting top $5$ images and apply $k=3$, using $3$ level of crops with a $0.2$ increment in scaling ratio i.e $S_{r}=[0.8,1,1.2]$. A stride of $s=40$ is used to ensure adequate overlap between frames in chosen datasets. For GroundedSAM [13], thresholds are set as follows: IoU-0.4, bounding box-0.25, and text-0.25. A padding of $px=20$ pixels delineates borders between instance masks. Overlap ratio evaluation uses a voxel size of $\epsilon=0.02$ and an overlap threshold of $\rho=0.3$. DBSCAN post-processing employs an epsilon of 0.1 and a minimum cluster size of 20 points. GPT-4 [6] is configured with a temperature of 0.
IV-B3 Filtering and Post-Processing
Large background objects (walls, ground, roof, ceiling) and objects with bounding boxes $>$ 95% of image area are excluded to prevent their feature vectors from exhibiting similarity to foreground objects, adversely affecting recall and score distribution. In DBSCAN post-processing, clusters with points $\geq$ 80% of the largest cluster are treated as separate instances with unique IDs and attributes. In cases where an object is undetectable by GPT-4 [6], instances are assigned RAM++ [35] names and given simplified captions: ”an object in a scene”.
IV-C Quantitative Evaluation
@@
The proposed method is evaluated using standard metrics: mean recall accuracy (mAcc), frequency-weighted IoU (F-mIoU), and average precision (AP) at IoU thresholds [0.5:0.05:0.95], along with AP50 and AP25 as defined in ScanNet [1]. For open vocabulary performance, similar to [27, 34], we retrieved 3D masks using ground truth labels with the prompt ‘an object in a scene’. We downsample retrived and ground truth masks to 0.25cm voxel size and applied a nearest neighbor search for intersecting points identification. We compared our results on the Replica [2] dataset with state-of-the-art models [27, 31, 11, 34], using identical prompts and foundation models.
IV-D Qualitative Evaluation
@@
Extensive qualitative assessment for open vocabulary instance retrieval, annotation, segmentation, and spatial reasoning was performed with manual human evaluation. For open vocabulary instance retrieval, over 1,000 queries regarding instances, affordances, properties, and relative queries were made. Performance was evaluated using CLIP [3] and four fusion schemes: Scheme 1 represents direct aggregation of multiscale (Eq. 4) and multiview features (Eq. 5), Scheme 2 represents updated multi-view features (Eq. 7), Scheme 3 represents updated multi-scale features (Eq. 6) with increased crop expansion ratios ($S_{r}=[1,2,4]$), and Scheme 4 represents combination of updated multi-view (Eq. 7) and multi-scale (Eq. 6) features fusion formulation.
The Annotation and Segmentation capabilities of the proposed approach were evaluated through manual verification of label assignments and mask merging. For spatial reasoning, 70 complex questions were administered across all scenes using a large language model (see Section III-F), assessing prompting strategy’s viability for spatial reasoning.
V Results and Discussion
V-A Ablation Studies
@@
To assess the influence of hyperparameters, ablation studies were conducted on Crop Level $k$, Top Images $m$, and Crop Ratios $S_{r}$ using quantitative metrics. Top Images $m$ affects multiview feature fusion (Eq. 7), representing the feature vectors for aggregation. Crop Ratio $S_{r}$ and Crop Level $k$ affect multi-scale feature fusion (Eq. 4), determining crop size and quantity for feature vector aggregation. Crop Level $k$ amplifies Crop Ratio $S_{r}$’s effect, as a higher $k$ with the same $S_{r}$ results in larger crops.
@@
Similar to OpenMask3D [27], we found that extreme values of these hyperparameters deteriorate results. A low $m$ reduces redundancy, while a high $m$ may include bad images, as shown in Table I. Lower values of $S_{r}$ and $k$ may not harm the model but could introduce redundancy. Larger $S_{r}$ values add context but may saturate similarity scores.
@@
Additionally, an ablation study with different CLIP model variants was also conducted. As shown in Table II, larger CLIP variants improved overall performance. As an example, Fig. 4 illustrates that a larger CLIP model better associates properties with the queried object, distinguishing between “single sofa” and “double sofa”.
CLIP Models Replica [2]
mAcc F-mIoU AP AP50 AP25
ViT-L-14 38.9 40.2 9.3 20.5 30.3
ViT-H-14 42.6 40.9 9.9 21.6 31.6
TABLE II: Ablation study of CLIP models. Influence of different CLIP Models on instance recall performance.
“double sofa”
ViT-L-14
ViT-H-14
“single sofa”
ViT-L-14
ViT-H-14
Figure 4: Text-Instance Similarity Heatmaps. For a text query, per-instance cosine similarity heatmaps for different CLIP models. $\blacksquare$: max, $\blacksquare$: min similarity.
V-B Open Vocabulary Instance Retrieval
V-B1 Quantitative Comparison with Baseline Methods
The proposed method demonstrates comparable or better performance than baselines on quantitative metrics, as shown in Tables III and IV. These compare segmentation mask accuracy and precision in response to open vocabulary queries against ground truth masks. We followed the original ConceptGraph [34] and OpenMask3D [27] setups for fair comparison. Overall, our method performed on par or better across all metrics and datasets.
@@ V-B1 Quantitative Comparison with Baseline Methods
TABLE III: Comparison of open-vocabulary segmentation results with ConceptGraph[34] setup.
@@
TABLE IV: Comparison of open-vocabulary segmentation results with OpenMask3D [27] setup.
V-B2 Qualitative Comparison with Baseline Methods
The quantitative evaluation was primarily designed for closed vocabulary assessments, relying on recall accuracy that does not reflect real-world requirements for open vocabulary queries. These methods, dependent on the quantity of mask proposals [31], may not accurately represent true performance.
To address these limitations, we provide comprehensive qualitative comparisons with baseline works in Fig. 5. The goal is to assess the ability to recall the correct segmentation mask for open vocabulary queries, assigning high similarity scores to relevant objects and lower scores to irrelevant ones. Specifically, our proposed method showed better 2D to 3D association and distribution of similarity scores with proposed multi-scale and multi-view feature fusion formulation (Eq. 6 and Eq. 7). As shown in Fig. 5, for the queries “a picture on wall” and “an empty vase”, both baseline methods recalled incorrect objects, while our method works perfectly.
@@ V-B2 Qualitative Comparison with Baseline Methods
Instance QueriesAffordance Queries
“owl”
“bottle case”
“place to sit and work”
“place to cook”
Property QueriesRelative Queries
“black trashcan”
“green towel”
“place to sit under display”
“green towel next to sink”
Figure 7: Text-Instance Similarity per Query Type. cosine similarity score of each instance for different query typers. $\blacksquare$: max, $\blacksquare$: min similarity.
V-B3 Assessment of Feature Fusion Schemes
We conducted a qualitative evaluation of feature fusion schemes as defined in Sec. IV-D. The results are summarized in Table V.
For Instance, Property, and Affordance queries, performance across all schemes was similar. However, for Relative queries, Scheme 2 and Scheme 3 with our proposed multi-scale and multi-view fusion formulations (Eqs. 7 and 6) outperformed Scheme 1. Scheme 4, incorporating both the proposed formulations, achieved the best recall accuracy for segmentation masks.
Similarity score heatmaps as shown in Fig. 8 show that Scheme 1 often wrongly assigned the highest score to the largest instance. In contrast, the updated fusion formulations in Schemes 2, 3, and 4 improved the recall of instance masks and similarity score distribution, with Scheme 4 performing the best overall.
@@ V-B3 Assessment of Feature Fusion Schemes
TABLE V: Evaluation of feature fusion schemes. Accuracy of fusion schemes for retrieval with “Inst.” (instance), “Aff.” (affordance), “Prop.” (property), and “Rel.” (relative) text queries, as assessed by a human evaluator.
@@
V-C Open Set Annotation and Segmentation
Annotation accuracy for directly assigned label $n$ with maximum prediction score $S^{\prime}_{\text{pred}}$ and LLM refined top $m$ labels corresponding to $S^{\prime}_{\text{pred}}$ label $n^{\prime}$ (discussed in III-C), were manually verified across all Replica [2] and ScanNet [1] scenes. To evaluate open set segmentation, under-merges and over-merges were counted and classified as faulty. Corresponding experiments in Table VI indicates that LLM labels $n^{\prime}$ are more accurate than direct labels $n$, highlighting that redundancy improves label (name) assignment while filtering undesired instances (Sec. V-A). This leads to a slight improvement in mask merging accuracy. Additionally, LLM labels $n^{\prime}$ are more concise than direct labels $n$.
@@ V-C Open Set Annotation and Segmentation
TABLE VI: Qualitative evaluation of segmentation and annotation accuracy. For Direct Label ($n$) and LLM Label ($n^{\prime}$), the annotation and merge accuracy of segmentation masks, as assessed by a human evaluator.
V-D Complex Spatial Queries
To assess spatial reasoning, we posed 70 complex spatial questions (Example, Fig. 1 “Which sink is closer to bed”) across scenes as specified in Sec. III-F. The manual assessment showed that, with our prompting strategy, LLM demonstrated effective reasoning in 3D space, over constructed representation. LLM exhibited higher accuracy in for scenes from Replica [2] (0.83), compared to ScanNet [1] (0.68). This decline in performance in larger scenes is attributed to a higher incidence of merging flaws and labels associated to larger scenes.
V-E Limitations
The effectiveness of this approach is constrained by the capabilities of its underlying foundation models and the occurrence of merging errors. The method leverages CLIP for image-text association and GroundedSAM for 2D mask generation, making its performance directly tied to the robustness of these foundational models. Furthermore, the method’s spatial reasoning and annotation accuracy depends on LLM [6] and are impacted by occasional merging faults.
VI CONCLUSION
In conclusion, this study presents a scalable and incremental framework for open world 3D scene understanding, addressing limitations of current non-incremental methods. By leveraging 2D foundation models, our approach constructs detailed instance-level 3D scene representations, efficiently tracking and associating instance-specific information such as feature vectors, names, and captions. The proposed feature fusion schemes enhance the model’s ability to contextualize and interpret complex queries. Additionally, the use of large language models for automatic annotation and advanced spatial reasoning tasks demonstrates the method’s versatility and robustness. Comprehensive evaluations show that our method achieves superior zero-shot generalization compared to state-of-the-art solutions. In future, we plan to explore spatio-temporal reasoning in 3D dynamic scenes and extend the method from indoors to large-scale outdoor environments.

### v2

@@ II-B2 Feature Fusion
The fusion schemes in Eq. 3 and Eq. 4, lack contextual information, resulting in suboptimal performance for contextual/relative queries (Table IV). [29] and our ablations (Sec. III-C1) indicate that while multiscale crops (important for encompassing contextual information) enhance accuracy, larger crops diminish overall object recall. To mitigate these issues, we propose a modified multiscale feature fusion scheme Eq. 5. This approach employs weighted aggregation of multiscale crop feature vectors, where $f_{1}$ represents the best-fit crop’s feature vector and $\varepsilon$ is a small number $\approx$ 1e-8.
$f_{\text{MS}}=\frac{1}{k}\sum_{i=1}^{k}\left(\frac{f_{1}\cdot f_{i}}{\max(\|f_{1}\|_{2}\cdot\|f_{i}\|_{2},\varepsilon)}\right)\cdot f_{i}$ (5)
For multiview feature integration, inspired by the per-pixel feature representation in [11], we propose incorporating a global feature vector $f_{\text{G}}$ from the entire image while synthesizing our per instance feature representation, defined as:
$f_{\text{MV}}=\frac{1}{m}\sum_{i=1}^{m}f_{\text{MS}_{i}}+\left(\frac{f_{\text{MS}_{i}}\cdot f_{\text{G}_{i}}}{\max(\|f_{\text{MS}_{i}}\|_{2}\cdot\|f_{\text{G}_{i}}\|_{2},\varepsilon)}\right)\cdot f_{\text{G}_{i}}$ (6)
II-C Instance Retrieval & Segmentation
Given a map $\mathcal{M}$, open-vocabulary 3D object search or 3D instance retrieval and segmentation is performed in two stages. First, a query $\mathcal{K}$ is processed using the CLIP [3] text encoder to obtain the feature vector $f_{\mathcal{K}}$. Second, cosine similarity scores $\{S_{\text{score}}\}$ are computed with all 3D instances. The segmentation mask of the 3D instance with the highest similarity score is retrieved as the most likely response to $\mathcal{K}$.
II-D Spatial Reasoning
For queries requiring complex spatial reasoning, our approach involves in-context learning, leveraging large context window of LLM [6], enabling complex spatial reasoning based on coherent 3D representations and associated metadata such as mask labels, centroids, bounding boxes and detailed caption available for the constructed scene $\mathcal{M}$. :=$\mathcal{M}^{\prime}:=\mathcal{M}\setminus\{\mathcal{P},f_{\text{MV}}\}$ is provided to LLM along with a system prompt, which is crafted using the following prompting strategy:
@@ II-D Spatial Reasoning
III Experiments
III-A Implementation Details
III-A1 Models Utilized
@@
GroundedSAM [13] (a method based on RAM++ [39] (ram_plus_swin_large_14m), GroundingDINO [40] (groundingdino_swint_ogc), and SAM [4] (sam_vit_h_4b8939)) is employed for generating instance segmentation masks and bounding boxes. GPT-4V [6] (gpt-4-vision-preview, gpt-4-1106-preview) is used for detailed captions & names and spatial reasoning. The CLIP encoder [3] (ViT-H-14 pre-trained on laion2b_s32b_b79k dataset) is used for instance feature vectors.
III-A2 Hyper parameter Settings
@@
Hyper-parameters, determined through ablation studies on Replica [2] (Sec: III-C1), are consistent across datasets. We set $m=5$ to select the top 5 images and use $k=3$ with 3 crop levels and a scaling ratio increment of 0.2 ($S_{r}=[0.8,1,1.2]$). A stride of $s=40$ ensures frame overlap. For GroundedSAM [13], thresholds are: IoU-0.4, bounding box-0.25, and text-0.25. Padding of $px=20$ pixels delineates instance mask borders. Overlap evaluation uses voxel size $\epsilon=0.02$ and threshold $\rho=0.3$. DBSCAN uses epsilon 0.1 and a minimum cluster size of 20 points. GPT-4 [6] is set with a temperature of 0.
III-A3 Filtering and Post-Processing
[…]
III-B Evaluation Details
III-B1 Datasets
Multiple scenes from the semi-synthetic dataset Replica [2] (room0, room1, room2, office0, office1, office2, office3, office4) and real-world dataset ScanNet [1] (scene0000_00, scene0034_00, scene0164_03, scene0525_01, scene0549_00) were used for thorough qualitative and quantitative evaluation. Similar to previous studies [37, 11], a limited number of scenes were selected due to the extensive manual human evaluation.
III-B2 Quantitative Evaluation
@@
The proposed method is evaluated using standard metrics: mean recall accuracy (mAcc), frequency-weighted IoU (F-mIoU), and average precision (AP) at IoU thresholds [0.5:0.05:0.95], including AP50 and AP25 from ScanNet [1]. For open vocabulary performance, as in [29, 37], 3D masks were retrieved with ground truth labels and the prompt: “an {object} in a scene”. Masks were downsampled to 0.25cm voxel size, followed by nearest neighbor search for intersecting points. Results on the Replica [2] dataset were compared against state-of-the-art models [29, 35, 11, 37], using identical prompts and foundation models and simple feature fusion formulation (Eq. 4 and Eq. 3).
III-B3 Qualitative Evaluation
@@
Extensive qualitative assessments for open vocabulary instance retrieval, annotation, segmentation, and spatial reasoning were performed through manual human evaluation. For open vocabulary instance retrieval, over 1,000 queries regarding instances (e.g. “musical instrument”), affordances (e.g., “place to sit and work”), properties (e.g., “green towel”), and relative queries (e.g., “green towel next to sink”) were made. Performance was evaluated based CLIP [3] based instance retrieval (see Sec.  II-C) for four feature fusion schemes: Scheme 1 represents direct aggregation of multiscale (Eq. 3) and multiview features (Eq. 4), Scheme 2 represents updated multiview features (Eq. 6), Scheme 3 represents updated multiscale features (Eq. 5) with increased crop expansion ratios ($S_{r}=[1,2,4]$), and Scheme 4 represents combination of updated multiview (Eq. 6) and multiscale (Eq. 5) features fusion formulation. The annotation and segmentation capabilities of the proposed approach were evaluated through manual verification of label assignments and mask merging. For spatial reasoning, 70 complex reasoning questions (e.g., “I’m feeling cold what should I do?”) were administered across all scenes using a large language model (see Sec.  II-D), assessing viability of proposed prompting strategy for spatial reasoning.
III-C Results and Discussion
III-C1 Ablation Studies
@@
To assess the influence of hyperparameters, using [29] experiment setup; ablation studies were conducted on Crop Level $k$, Top Images $m$, and Crop Ratios $S_{r}$ using quantitative metrics. Top Images $m$ affects multiview feature fusion (Eq. 6), representing the feature vectors for aggregation. Crop Ratio $S_{r}$ and Crop Level $k$ affect multiscale feature fusion (Eq. 3), determining crop size and quantity for feature vector aggregation. Crop Level $k$ amplifies Crop Ratio $S_{r}$’s effect, as a higher $k$ with the same $S_{r}$ results in larger crops.
@@
Similar to [29], we found that extreme values of these hyperparameters deteriorate results. A low $m$ reduces redundancy, while a high $m$ may include bad images, as shown in Table I. Lower values of $S_{r}$ and $k$ may not harm the model but could introduce redundancy. Larger $S_{r}$ values add context, but can saturate the distribution of similarity scores.
@@
III-C2 Quantitative Comparison with Baseline Methods
The proposed method demonstrates comparable or better performance than baselines on quantitative metrics, as shown in Tables II and III. These tables compare segmentation mask accuracy and precision in response to open vocabulary queries against ground truth masks. We followed the original ConceptGraph [37] and OpenMask3D [29] setups for fair assessment. Overall, our method performed on par or better across all metrics and datasets.
@@ III-C2 Quantitative Comparison with Baseline Methods
TABLE II: Comparison of open-vocabulary segmentation results with ConceptGraph[37] setup.
@@
TABLE III: Comparison of open-vocabulary segmentation results with OpenMask3D [29] setup.
III-C3 Qualitative Comparison with Baseline Methods
The quantitative evaluation was primarily designed for closed vocabulary assessments, relying on recall accuracy that does not reflect real-world requirements for open vocabulary queries. Additionally, these evaluations, dependent on the quantity of mask proposals [35], might not accurately represent true performance against closed set ground truth masks.
To address these limitations, we provide comprehensive qualitative comparisons with baseline works in Fig. 4. The goal is to assess the ability to recall the correct segmentation mask for open vocabulary queries, assigning high similarity scores to relevant objects and lower scores to irrelevant ones. Specifically, our proposed method showed better 2D to 3D association and distribution of similarity scores with proposed multiscale and multiview feature fusion formulation (Eq. 5 and Eq. 6). As shown in Fig. 4, for the queries “a picture on wall” and “an empty vase”, both baseline methods recalled incorrect objects, while our method works perfectly.
@@ III-C3 Qualitative Comparison with Baseline Methods
III-C4 Assessment of Feature Fusion Schemes
We conducted a qualitative evaluation of feature fusion schemes as defined in Sec. III-B3. The results are summarized in Table IV.
For instance, property, and affordance queries, performance across all schemes was similar. However, for relative queries, Scheme 2 and Scheme 3 with our proposed multiscale and multiview fusion formulations (Eqs. 6 and 5) outperformed Scheme 1. Scheme 4, incorporating both of the proposed formulations, achieved the best recall accuracy. Additionally, similarity score heatmaps as shown in Fig. 5 shows that Scheme 1 often wrongly assigned the highest score to the largest instance. In contrast, the updated fusion formulations in Schemes 2, 3, and 4 improved the recall of instance masks and similarity score distribution, with Scheme 4 performing the best overall.
@@ III-C4 Assessment of Feature Fusion Schemes
TABLE IV: Evaluation of feature fusion schemes. Accuracy of fusion schemes for retrieval with “Inst.” (instance), “Aff.” (affordance), “Prop.” (property), and “Rel.” (relative) text queries, as assessed by a human evaluator.
@@
III-C5 Open Set Annotation and Segmentation
Annotation accuracies for directly assigned labels $n$ using the maximum prediction score $S^{\prime}_{\text{pred}}$ and labels $n^{\prime}$, where top $m$ labels based on $S^{\prime}_{\text{pred}}$ are refined by LLM (see Sec. II-B1), were manually verified across all Replica [2] and ScanNet [1] scenes. To evaluate open set segmentation, mask merging accuracy is determined by counting and classifying under-merges and over-merges as faulty. Evaluations summarized in Table V show that LLM labels $n^{\prime}$ are more accurate than direct labels $n$. Additionally, this helps to filter out undesired large instances (see Sec. III-A3), resulting in slight improvement in mask merging accuracy. Furthermore, we found that LLM labels $n^{\prime}$ are more concise than direct labels $n$.
@@ III-C5 Open Set Annotation and Segmentation
TABLE V: Qualitative evaluation of segmentation and annotation accuracy. For Direct Label ($n$) and LLM Label ($n^{\prime}$), the annotation and merge accuracy of segmentation masks, as assessed by a human evaluator.
III-C6 Complex Spatial Reasoning
To assess spatial reasoning, as specified in Sec. III-B3, we posed complex spatial reasoning questions (Example, Fig. 1 “Which sink is closer to bed”) across all scenes. The manual assessment showed that, with our approach (see II-D), the LLM [6] demonstrated effective reasoning ability in 3D space over constructed representation. LLM exhibited higher accuracy in scenes from Replica [2] (0.83), compared to ScanNet [1] (0.68). This decline in performance can be attributed to a comparatively higher incidence of flaws in merging and label assignment in larger ScanNet [1] scenes.
III-D Limitations
The effectiveness of this approach is constrained by the capabilities of its underlying foundation models and the occurrence of merging errors. Utilizing CLIP [3] for image-text association and  [13] for 2D mask generation, performance correlates with these model’s robustness. Moreover, annotation accuracy and spatial reasoning depends on LLM’s [6] ability and context length, and are susceptible to occasional merging faults.
IV CONCLUSION
In conclusion, this study presents a scalable and incremental framework for constructing open set 3D scene representation for open world 3D scene understanding tasks, addressing the limitations of current methods. By leveraging 2D foundation models, our approach constructs detailed instance-level 3D scene representations, efficiently tracking and associating instance-specific information such as feature vectors, names, and captions. The proposed feature fusion schemes encompass contextual information, enhancing performance on relative queries. Additionally, the use of large language models facilitates robust automatic annotation and enables complex spatial reasoning over 3D scenes. Comprehensive evaluations show that our method achieves superior zero-shot generalization compared to state-of-the-art solutions. In the future, we plan to explore spatio-temporal reasoning in 3D dynamic scenes and extend the method from indoors to large-scale outdoor environments.